# What Did My Car Say? Autonomous Vehicle Explanation Errors, Context, and Personal Traits Impact Comfort, Reliance, Satisfaction, and Driving Confidence

### v1

@@ 4.3.2. Correlations of Initial Trust and Subjective Expertise with Outcomes
4.4. Self-Reported Rationale For Trust and Reliance Decisions
@@ 4.4.1. Factors Contributing to Rating Decisions
We explicitly inquired upon the basis of reliance decisions by asking participants to rate the relative importance of several factors on their reliance ratings (Figure 4). This was done after the rating task concluded. As a whole, we found that the accuracy of the explanation on ‘what’ the AV is doing was the most important consideration, followed by the explanation accuracy of ‘why’ the AV is doing it, the harm of the situation, the AV’s driving ability, and the difficulty of the driving situation. Prior scenes viewed and prior knowledge did not have much reported impact, though this is unsurprising, as these implicit judgments may be difficult for participants to self-describe. Many of these differences were significant based on ANOVA. Notably, harm was significantly more important than difficulty, while explanation accuracy (both ‘what’ and ‘why’ components) were significantly more important than AV driving ability. This latter effect may be in part due to the salience of these factors based on the study’s manipulation. We did not see significant differences between the importance of ‘what’ and ‘why’ accuracy.
Figure 4. Relative Importance of Factors on Reliance Decisions. This shows the mean importance rating of each factor with respect to a person’s reliance decision. Results show many of the factors of interest in this study were of high importance.
[…]
@@ 4.4.2. Feedback from Participants
Regarding exposure to explanation errors, one participant commented that they were “less confident in the reliability of [autonomous] vehicles” after exposure to errors, while another commented that when explanations “were wrong, [it made] my confidence in the system shaky.” These comments reiterate our general study findings on the negative impact of errors on reliance, as well as the holistic judgment of an AV’s ability in general. Some participants broke their decision-making down in terms of individual factor priority. For instance, “if the AV described the action it was taking incorrectly I was a lot less confident and consistently chose to take control myself … if the reason was incorrect, it still impacted my confidence level but not as much.” This reflects the general trend found in our individual factor analysis suggesting ‘what’ information may be more important than ‘why’ rationale, even if this trend was not found to be statically significant.
Regarding the impact of context, a participant noted, “I chose to take control myself in higher risk situations even if the AV did a good job navigating.” This supports this study’s overall findings on the importance of contextual factors like harm on decision-making.
@@ 5. Discussion
Autonomous vehicle (AV) explanation errors had a detrimental effect on all outcomes. Our results suggest that errors significantly reduced participant comfort relying on an AV, preference to rely on the AV instead of taking control themselves, and satisfaction with the explanation. These effects are unsurprising, as we would expect outcomes related to a person’s reliance decision or satisfaction with an explanation to reflect the system’s performance (Hoff and Bashir, 2015).
@@
We were surprised, however, to find crossover effects between the explanatory performance of the AV and a person’s confidence in the AV’s driving ability, given the AV’s actual demonstrated driving performance remained consistent across all conditions. This crossover effect alludes to the mental model of potential AV users, where evaluation of the explanatory performance of the vehicle and the driving performance of the vehicle are connected. Though connecting explanatory and driving performance may be an intuitive and correct assumption, empirically observing this crossover provides evidence supporting the importance of high quality explanatory interfaces, particularly within safety-critical systems. For the case of AVs, this means that – even if the AV’s driving performance is perfect – if the explanations produced by the AV are not accurate as well, people may still refuse to adopt AV technology. This insight likely generalizes to explanatory communications for other AI-based systems.
@@
The negative impact of errors increased with error magnitude and the potential harm of the error. Expanding on the effect of errors observed in our study, we find that the impact was commensurate with the magnitude of the errors presented. In the study presented here, ‘what’ and ’why’ errors (high condition) in combination had worse outcomes than ‘why’-only errors (low condition). On the surface, this is unsurprising, as more errors demonstrates lower system performance, and these results could simply be the cumulative effects of seeing errors on two parts of the explanation instead of just one.
@@
There remains a high probability that ‘what’ and ‘why’ information errors play distinct roles in the evaluation of the system’s performance, however. This would align with past work showing how description of action and description of justification may differentially impact the way a person interacts with autonomous systems (Koo et al., 2015; Kaufman et al., 2024a; Miller, 2019). When asked explicitly about the factors that contributed to their reliance decisions, participants reported that explanation accuracy (both ‘what’ and ‘why’ components) were both important. Qualitative assessment of participant feedback, however, suggests that ‘what’ information may be more important for reliance outcomes than ‘why’ rationale. This may be due to the potential harm that improper action (‘what’) can have on driving safety, in contrast to improper justification (‘why’). Concretely, a car incorrectly turning right into traffic can cause physical harm, but incorrectly justifying why it is turning right can not. We note that conclusively differentiating the impact of ‘what’ from ‘why’ errors independently will be left for future work.
@@
Even with accurate explanations, Comfort, Reliance Preference, and Confidence in the AV’s driving ability were fairly low. In the present study, we were primarily interested in the relative differences in outcome ratings based on errors, contextual factors, or individual differences like trust or expertise level. We note, however, that even in cases when the explanation presented was accurate, comfort, reliance, and confidence scores still hovered near the middle of the provided scale. Though the middling scores may be reflective of the overall challenging nature of the driving situations presented in the study, even the most straightforward of the scenarios presented did not have a reliance preference score greater than mid-range. These results reflect the trend found in a plethora of prior work supporting the premise that people do not generally trust or wish to adopt autonomous vehicles.
Contextual factors like the difficulty of driving in a specific situation and the perceived harm of that situation directly affect how people think, feel, and prefer to behave with AVs. We found evidence that driving difficulty and perceived harm directly influenced outcomes regardless of error condition. Specifically, higher contextual harm was associated with lower comfort, reliance, and confidence ratings, while higher difficulty was associated with higher reliance ratings. We did not find direct associations between harm or difficulty and explanation satisfaction.
@@
Explicitly, participants reported contextual harm as significantly more important than driving difficulty for their reliance decision. From a theoretical lens, a person’s concern with difficult driving may stem from the increased likelihood of harm in challenging situations, particularly when they may have concerns with an AV’s ability to perform. This leads us to hypothesize that an overall evaluation of potential for harm is the underlying factor influencing reliance decisions overall. By this lens, difficulty and AV driving performance matter because harm matters. Thus, we hypothesize that harm itself may be the underlying driver of reliance decisions overall.
@@
Contextual factors moderate the relationship between errors and outcomes in complex, sometimes counter-intuitive ways. By examining the interaction between contextual characteristics and the effects of error conditions, we found that – in some cases – the driving context may influence the amount of impact an error has. The general trend we observed for comfort relying on the AV, reliance preference, and confidence in the AV’s driving ability was that when difficulty increased, outcome scores decreased more in the low error condition compared to the accurate condition, with no differences seen between the accurate and high conditions. The same but opposite effect was found for harm: higher harm increased outcomes effect in the low condition more than the accurate condition, with no difference between the accurate and high condition.
@@
Trust and expertise influenced how people think, feel, and behave towards the AV. As discussed, how a person thinks and behaves towards AVs is not just a function of their external environment. We find evidence that internal characteristics, such as a participant’s trust and expertise, may also have played a role in the study’s results. Participants with higher initial trust or expertise tended to rate study outcomes higher on average. We also found that those with higher AV expertise trusted AVs more in general. As with the function on driving reliance described previously, these findings also align with Hoff and Bashir’s model of trust with AI systems, where a person’s prior experience impacts their reliance with the system (Hoff and Bashir, 2015). Those with higher expertise may also have other underlying characteristics, like an affinity for new technologies, which may have influenced their trust. We attribute the finding that priors were not reported as explicitly important for reliance decisions in our post-evaluation to the common difficulty articulating the impact of implicit dispositions.
@@ 5.1. Implications For Autonomous Vehicle Design and Research
Our results have important implications for future AV design and research. Understanding the consequences of AV explanation errors and contextual characteristics like driving difficulty and perceived harm are necessary first steps towards designing vehicles which may be more trustworthy, reliable, and satisfactory for people interacting with the AV. In this section, we will discuss how our study’s insights can be used to design more trustworthy AVs.
@@
Another major implication of this work is to orient AV designers to the potentially different outcomes that may result from interactions by people with different traits or while operating in driving environments. Specifically, context-aware or personalized explanations may be necessary in cases where the driving difficulty or perceived harm are classified as greater, or for people with little expertise or initial trust. It is likely that designers will need to focus on limitations on cognitive resources like attention and load, particularly for high difficulty scenarios. Other personal traits like risk preferences or personality may also be a basis for personalization (Böckle et al., 2021). Segmenting populations or isolating particular driving conditions to test future explanation designs would be an effective method to figure out how to meet the differing needs of diverse user groups (Kaufman et al., 2024b).
@@ Acknowledgements.
We would like to thank S. Sapre, R. Bhide, C. Lee, J. Molina, and E. Lee for simulator development and testing.

### v3

@@ 4.3.2. Correlations of Initial Trust and Subjective Expertise with Outcomes
4.4. Self-Reported Rationale for Reliance Decisions
@@ 4.4.1. Factors Contributing to Rating Decisions
We explicitly inquired upon the basis of reliance decisions by asking participants to rate the relative importance of several factors on their reliance ratings (Figure 4). This analysis helps clarify which elements participants explicitly prioritized, offering valuable guidance for designing explanations that align with user preferences and expectations. Ratings were given after the main scenario experiment concluded. As a whole, we found that the accuracy of the explanation on ‘what’ the AV is doing was the most important consideration (mean = 3.96, SE = 0.07), followed by the explanation accuracy of ‘why’ the AV is doing it (mean = 3.91, SE = 0.07), the harm of the situation (mean = 3.75, SE = 0.07), the AV’s driving ability (mean = 3.49, SE = 0.07), and the difficulty of the driving situation (mean = 3.42, SE = 0.06). Prior scenes viewed (mean = 2.71, SE = 0.08) and prior knowledge (mean = 2.61, SE = 0.08) did not have much reported impact, though this is unsurprising, as these implicit judgments may be difficult for participants to self-describe. Many factor differences were significant based on Kruskal-Wallis test ($\chi^{2}$ = 285.7, df = 6, p ¡ 0.001). For pairwise comparisons, we use Dunn’s test and adjust p-values using Bonferroni’s method. Notably, we find harm was significantly more important than difficulty (p ¡ 0.05), while explanation accuracy (both ‘what’ and ‘why’ components) were significantly more important than AV driving ability (both p ¡ 0.001). This latter effect may be in part due to the salience of these factors based on the study’s manipulation. We did not see significant differences between the importance of ‘what’ and ‘why’ accuracy.
@@ 4.4.2. Feedback from Participants
Regarding exposure to explanation errors, one participant commented that they were “less confident in the reliability of [autonomous] vehicles” after exposure to errors, while another commented that when explanations “were wrong, [it made] my confidence in the system shaky.” These comments reiterate our findings on the negative impact of errors on reliance, as well as the holistic judgment of an AV’s ability in general. Some participants broke their decision-making down in terms of individual factor priority. For instance, “if the AV described the action it was taking incorrectly I was a lot less confident and consistently chose to take control myself … if the reason was incorrect, it still impacted my confidence level but not as much.” This reflects the general trend found in our individual factor analysis suggesting ‘what’ information may be more important than ‘why’ rationale, even if this trend was not found to be statically significant.
Regarding the impact of context, a participant noted, “I chose to take control myself in higher risk situations, even if the AV did a good job navigating.” This supports this study’s overall findings on the importance of contextual factors like harm on decision-making.
@@ 5. Discussion
Autonomous vehicle (AV) explanation errors had a detrimental effect on all outcomes (RQ1). Our results suggest that errors significantly reduced participant comfort relying on an AV, preference to rely on the AV instead of taking control themselves, and satisfaction with the explanation. These effects are unsurprising, as we would expect a person’s reliance decision or satisfaction with an explanation to reflect the system’s performance (Hoff and Bashir, 2015).
@@
We were surprised, however, to find crossover effects between the explanatory performance of the AV and a person’s confidence in the AV’s driving ability, given the AV’s actual demonstrated driving performance remained consistent across all conditions. This crossover effect alludes to the mental model of potential AV users, where evaluation of the explanatory performance of the vehicle and the driving performance of the vehicle are connected. Though connecting explanatory and driving performance may be an intuitive and correct assumption, empirically observing this crossover provides evidence supporting the importance of high quality explanatory interfaces, particularly within safety-critical systems. Further highlighting this point, the finding that explanation accuracy was explicitly rated as more important than driving ability for reliance decisions underscores the role of transparency and comprehensibility in user acceptance. This highlights the need for AV systems to prioritize clear and accurate explanations of actions (‘what’) and rationales (‘why’) in building trust, and suggests that users may evaluate AV systems as much by their communications as by their actual driving performance. Taken together, in the case of AVs, this means that – even if the AV’s driving performance is perfect – if the explanations produced by the AV are not accurate as well, people may still refuse to adopt AV technology. This insight likely generalizes to explanatory communications for other AI-based systems.
Even with accurate explanations, Comfort, Reliance Preference, and Confidence in the AV’s driving ability were nowhere near ceiling level – hovering near the middle of the provided scale. Though the middling scores may be reflective of the overall challenging nature of the driving situations presented in the study, even the most straightforward scenarios did not have a mean reliance preference score greater than mid-range. These results reflect the trend found in a plethora of prior work supporting the premise that people do not generally trust or wish to adopt autonomous vehicles, reemphasizing this as an important area of continued study.
@@
The negative impact of errors increased with error magnitude and the potential harm of the error (RQ1 continued). Expanding on the effect of errors observed in our study, we find that the impact was commensurate with the magnitude of the errors presented. In the study presented here, ‘what’ and ‘why’ errors (high condition) in combination had worse outcomes than ‘why’-only errors (low condition). On the surface, this is unsurprising, as more errors demonstrates lower system performance, and these results could simply be the cumulative effects of seeing errors on two parts of the explanation instead of just one.
@@
There remains a high probability that ‘what’ and ‘why’ information errors play distinct roles in the evaluation of the system’s performance, however. This would align with past work showing how description of action and description of justification may differentially impact the way a person interacts with autonomous systems (Koo et al., 2015; Kaufman et al., 2024a; Miller, 2019). When asked explicitly about the factors that contributed to their reliance decisions, participants reported that explanation accuracy (both ‘what’ and ‘why’ components) were both important. Qualitative assessment of participant feedback, however, suggests that ‘what’ information may be more important for reliance outcomes than ‘why’ rationale. This may be due to the potential harm that improper action (‘what’) can have on driving safety, in contrast to improper justification (‘why’). Concretely, a car incorrectly turning right into traffic can cause physical harm, but incorrectly justifying why it is turning right can not. This finding highlights that explanation accuracy, particularly for ‘what’ information, is crucial not only for user comfort but also for the perception of the AV’s reliability and overall system competence. We note that conclusively differentiating the impact of ‘what’ from ‘why’ errors independently will be left for future work.
@@
Contextual factors like the difficulty of driving in a specific situation and the perceived harm of that situation directly affect how people think, feel, and prefer to behave with AVs (RQ2). We found evidence that driving difficulty and perceived harm directly influenced outcomes regardless of error condition. Specifically, higher harm was associated with lower comfort, reliance, and confidence ratings, while higher difficulty was associated with higher reliance ratings. We did not find direct associations between harm or difficulty and explanation satisfaction.
@@
Explicitly, participants reported contextual harm as significantly more important than driving difficulty for their reliance decision. Theoretically, this may be because difficult driving situations often entail a higher risk of harm, especially when participants are uncertain about the AV’s ability to handle these challenges. We hypothesize that reliance decisions are grounded in an overall evaluation of potential harm, with driving difficulty acting as one indicator of this risk. Accordingly, difficulty and AV driving performance matter because harm matters, making harm itself the underlying driver of reliance decisions.
@@
Contextual factors moderate the relationship between errors and outcomes in complex, sometimes counter-intuitive ways (RQ2 continued). By examining the interaction between contextual characteristics and the effects of error conditions, we found that – in some cases – the driving context may influence the amount of impact an error has. The general trend we observed for comfort relying on the AV, reliance preference, and confidence in the AV’s driving ability was that when difficulty increased, outcome scores decreased more in the low error condition compared to the accurate condition, with no differences seen between the accurate and high conditions. The same but opposite effect was found for harm: higher harm increased outcomes effect in the low condition more than the accurate condition, with no difference between the accurate and high condition.
@@
Trust and expertise influenced how people think, feel, and behave towards the AV (RQ3). As discussed, how a person thinks and behaves towards AVs is not just a function of their external environment. We find evidence that internal characteristics, such as a participant’s trust and expertise, may also have played a role in the study’s results. Participants with higher initial trust or expertise tended to rate study outcomes higher on average. We also found that those with higher AV expertise trusted AVs more in general. As with the function on driving reliance described previously, these findings also align with Hoff and Bashir’s model of trust with AI systems, where a person’s prior experience impacts their reliance with the system (Hoff and Bashir, 2015). Those with higher expertise may also have other underlying characteristics, like an affinity for new technologies, which may have influenced their trust. We attribute the finding that priors were not reported as explicitly important for reliance decisions in our post-evaluation to the common difficulty articulating the impact of implicit dispositions.
@@ 5.1. Implications For Autonomous Vehicle Design and Research
Our results have important implications for future AV design and research (RQ4). Understanding the consequences of AV explanation errors and contextual characteristics like driving difficulty and perceived harm are necessary first steps towards designing vehicles which may be more trustworthy, reliable, and satisfactory for people interacting with the AV. By using current, multimodal XAI methods of explanation presentation – visual and auditory cues of both ‘what’ and ‘why’ information – this study’s findings can directly apply to current AV explanation design. In this section, we will discuss how our study’s insights can be applied to build more trustworthy AVs.
@@
Another major implication of this work is to orient AV designers to the potentially different outcomes that may result from interactions by people with different traits or while operating in driving environments. In AVs, user familiarity with driving norms allows for immediate identification of errors, which intensifies the effect on trust and reliance. This contrasts with fields where users cannot as easily detect errors without domain expertise, highlighting the importance of context-specific trust calibration mechanisms for XAI systems. Specific to the case studied here, context-aware or personalized explanations may be necessary in cases where the driving difficulty or perceived harm are classified as greater, or for people with little expertise or initial trust. It is likely that designers will need to focus on limitations on cognitive resources like attention and load, particularly for high difficulty scenarios. In high-risk domains like AVs, these factors may be particularly important, as the potential consequences of explanation errors are immediate and tangible. We thus contribute to an emerging understanding of how explanation accuracy and error detectability shape user perceptions in safety-critical environments. Other personal traits like risk preferences or personality may also be a basis for personalization (Böckle et al., 2021). Segmenting populations or isolating particular driving conditions to test future explanation designs would be an effective method to figure out how to meet the differing needs of diverse user groups (Kaufman et al., 2024b).
@@ Acknowledgements.
We would like to thank Saumitra Sapre, Rohan Bhide, Chloe Lee, Janzen Molina, and Emi Lee for driving scenario development and simulator testing.